# A topological constraint on clean Lagrangian intersections from $\mathbb{Q}$-valued augmentations

### v1

@@ Abstract
For a knot $K$ in $\mathbb{R}^{3}$, consider its conormal bundle $L_{K}$ in $T^{*}\mathbb{R}^{3}$. We prove that if $K$ is the connected sum $K^{\prime}\#3_{1}$ of an arbitrary knot $K^{\prime}$ and the trefoil knot $3_{1}$ (or its mirror), then there is no compactly supported Hamiltonian diffeomorphism $\varphi$ on $T^{*}\mathbb{R}^{3}$ such that $\varphi(L_{K})$ and the zero section $\mathbb{R}^{3}$ intersect cleanly along the unknot in $\mathbb{R}^{3}$. For the proof, we utilize the knot DGA, whose homology was proved to be isomorphic to the knot contact homology by Ekholm, Etnyre, Ng and Sullivan, with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$. We discover a certain algebraic constraint imposed on any augmentation of the knot DGA of $K=K^{\prime}\#3_{1}$ to a field $\mathbb{k}$. This constraint is non-trivial for $\mathbb{k}=\mathbb{Q}$ and plays a key role in the proof, but trivial if $\mathbb{k}$ is quadratically closed or if an augmentation factors through the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ obtained by setting $U=1$.
@@ 1 Introduction
It is a Lagrangian submanifold of $T^{*}\mathbb{R}^{3}$ and intersects $\mathbb{R}^{3}$ cleanly along the knot $K$. Let $\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ denote the set of Hamiltonian diffeomorphisms on $T^{*}\mathbb{R}^{3}$ associated to compactly supported Hamiltonians on $T^{*}\mathbb{R}^{3}$. Using these notations, we ask a question which was observed by the author in [17].
Question 1.1.
@@
We remark that it is essential that $\varphi$ is a Hamiltonian diffeomorphism. Indeed, for any pair of knots $(K_{0},K_{1})$ in $\mathbb{R}^{3}$, there exists a $C^{\infty}$ isotopy $(\varphi^{t})_{t\in[0,1]}$ on $T^{*}\mathbb{R}^{3}$ with compact support such that $\varphi^{1}(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$. See [17, Proposition 4.1].
@@
As a special case, the answer to Question 1.1 when $K$ is the unknot is ‘no’. This can be deduced from the results by Ganatra-Pomerleano [10, Proposition 6.29], Smith-Wemyss [21, Proposition 4.10] or Asplund-Li [1, Theorem 1.3] about clean Lagrangian intersections in the plumbings of two copies of the cotangent bundle of $S^{3}$ along the unknots. These results are concerned with Smith’s question on the ‘persistence (or rigidity) of unknottedness of Lagrangian intersections’ [10, Question 1.5]. For its relation to Question 1.1, see [17, Section 1].
@@
In a general case, under the assumption that $\varphi(L_{K})$ for $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ and $\mathbb{R}^{3}$ intersect cleanly along a knot in $\mathbb{R}^{3}$, the author proved [17, Theorem 1.1] by using the (framed) knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$, which is a differential graded algebra (DGA) defined by Ng [14]. [17, Theorem 1.1] is stated by using the augmentation variety in $(\mathbb{C}^{*})^{2}$ derived from the set of augmentations of the knot DGA to $\mathbb{C}$. This theorem leads us to the same answer to Question 1.1 as above when $K$ is the unknot [17, Theorem 1.3] and we can also deduce some constraints on the knot type of $\varphi(L_{K})\cap\mathbb{R}^{3}$ when $K$ has certain non-trivial knot types [17, Theorem 1.4, 1.5]. However, we cannot deduce any constraint from [17, Theorem 1.1] on the knot type of $K$ if we suppose that $\varphi(L_{K})\cap\mathbb{R}^{3}$ is the unknot in $\mathbb{R}^{3}$. For the reason, see Remark 1.5 below.
Main results. Let $3_{1}$ denote the right-handed trefoil knot in $\mathbb{R}^{3}$ and $\overline{3_{1}}$ denote its mirror. The goal of this paper is to prove the next theorem which cannot be deduced from [17, Theorem 1.1].
Theorem 1.1 (Theorem 3.10).
@@
Let $K$ be the connected sum of an arbitrary knot in $\mathbb{R}^{3}$ with $3_{1}$ or $\overline{3_{1}}$. Then, there is no $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K})$ intersects $\mathbb{R}^{3}$ cleanly along the unknot in $\mathbb{R}^{3}$.
@@
For the proof, the geometric arguments in the present paper are similar to those in [17], which are outlined as follows: Let $K_{0}$ and $K_{1}$ be knots in $\mathbb{R}^{3}$. For $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$, suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ intersect cleanly along $K_{1}$. We use the results by Ekholm-Etnyre-Ng-Sullivan [7, 4] which show that the Chekanov-Eliashberg DGA of the unit conormal bundle of $K_{i}$ is stable tame isomorphic to the knot DGA of $K_{i}$ for $i=0,1$. The knot DGA, introduced by Ng [12, 13], is defined combinatorially from a braid representation of a knot. We then apply a DGA map associated to an exact Lagrangian cobordism between the unit conormal bundles of $K_{0}$ and $K_{1}$. This is a part of the symplectic field theory introduced by Eliashberg-Givental-Hofer [9]. Consequently, we obtain a DGA map from the knot DGA of $K_{0}$ to that of $K_{1}$. (Precisely, we need a base change on the coefficient ring of a DGA.)
The feature of the present paper is that we utilize the knot DGA, which we denote here by $(\mathcal{A}^{\mathrm{knot}}_{*}(K),\partial_{K})$ for an oriented knot $K$, with coefficients in the ring
$\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]\cong\mathbb{Z}[H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})].$
Here, $S^{*}\mathbb{R}^{3}$ denotes the unit cotangent bundle of $\mathbb{R}^{3}$ and $\Lambda_{K}$ denotes the unit conormal bundle of $K$. For details, we refer to [15, 4, 16]. We note that in [17], referring to [14, 7], the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]\cong\mathbb{Z}[H_{1}(\Lambda_{K})]$ was used.
To prove Theorem 1.1, we show the following algebraic property of the knot DGA under an assumption on the isotopy class of $K$. Let $K^{\prime}\#3_{1}$ denote the connected sum of a knot $K^{\prime}$ in $\mathbb{R}^{3}$ and the trefoil knot $3_{1}$.
Lemma 1.2 (Lemma 3.9).
Let $\mathbb{k}$ be a field and $K^{\prime}$ be a knot. Then, for any augmentation $\varepsilon\colon\mathcal{A}_{*}^{\mathrm{knot}}(K^{\prime}\#3_{1})\to\mathbb{k}$ of the knot DGA of $K^{\prime}\#3_{1}$ to $\mathbb{k}$,
$D_{\varepsilon}\coloneqq(\varepsilon(\mu)+1-\varepsilon(\mu)^{-1}\varepsilon(U))^{2}-4(\varepsilon(\mu)-1)$
satisfies $D_{\varepsilon}\in\{k^{2}\mid k\in\mathbb{k}\}$.
Here, $\varepsilon(\mu)$ and $\varepsilon(U)$ are abbreviations of $\varepsilon(\mu\cdot\mathbb{1})$ and $\varepsilon(U\cdot\mathbb{1})$ for the unit $\mathbb{1}$ of the knot DGA. In fact, $D_{\varepsilon}$ is the discriminant of a quadratic equation in $\mathbb{k}$ determined by $\varepsilon$ (see the proof of Lemma 3.9). Note that if $\varepsilon$ factors through the knot DGA over $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ obtained by setting $U=1$, then $\varepsilon(U)=1$ and we have an identity
$(\varepsilon(\mu)+1-\varepsilon(\mu)^{-1})^{2}-4(\varepsilon(\mu)-1)=(\varepsilon(\mu)-1+\varepsilon(\mu)^{-1})^{2},$
so the constraint on $D_{\varepsilon}$ is trivial.
For the proof of Theorem 1.1, we use the constraint from Lemma 1.2 in the case of $\mathbb{k}=\mathbb{Q}$, that is, $D_{\varepsilon}$ is the square of some rational number for any augmentation $\varepsilon$ of the knot DGA of $K^{\prime}\#3_{1}$ to $\mathbb{Q}$. (In fact, for this purpose, we may choose other fields such as $\mathbb{Q}[\sqrt{-1}]$ and $\mathbb{F}_{5}$. See Remark 3.11.) It is crucial that $\mathbb{Q}$ is not quadratically (and algebraically) closed, and we remark that the constraint on $D_{\varepsilon}$ is trivial for any $\varepsilon$ if $\mathbb{k}=\mathbb{C}$.
@@
In order to study the knot DGA of a knot $K$ (or the Chekanov-Eliashberg DGA of $\Lambda_{K}$), the set of augmentations to $\mathbb{k}$ has been used with $\mathbb{k}=\mathbb{C}$ or a finite field $\mathbb{F}_{p}$ for a prime number $p$. For examples of applications, see Ng [14, Section 5.2 and 5.3] and Cornwell [2] about the augmentation polynomial when $\mathbb{k}=\mathbb{C}$ and [14, Section 5.1] about the augmentation number when $\mathbb{k}=\mathbb{F}_{p}$. On the other hand, the rational numbers $D_{\varepsilon}$ from Lemma 1.2 when $\mathbb{k}=\mathbb{Q}$ are derived in a way different from those previous researches.
Lastly, we provide a theorem which is a refinement of [17, Theorem 1.1]. For an oriented knot $K$ in $\mathbb{R}^{3}$, let $V_{K}$ denote the augmentation variety defined by [16, Definition 5.1] as a subset of $(\mathbb{C}^{*})^{3}$. Let $(x,y,Z)$ be the coordinate of $\mathbb{C}^{3}$.
@@
For oriented knots $K_{0}$ and $K_{1}$ in $\mathbb{R}^{3}$, suppose that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ intersects $\mathbb{R}^{3}$ cleanly along $K_{1}$. Then, there exist integers $b,c$ and $d$ for which we have an injective map
$V_{K_{1}}\hookrightarrow V_{K_{0}}\colon(x,y,Z)\mapsto(xy^{b}Z^{c},yZ^{d},Z).$
@@
Identifying $\{(x,y,Z)\in(\mathbb{C}^{*})^{3}\mid Z=1\}$ with $(\mathbb{C}^{*})^{2}$, the augmentation variety considered in [17] coincides with $V_{K}\cap\{Z=1\}\subset(\mathbb{C}^{*})^{2}$. It follows from Theorem 1.4 that there exists an injective map
$V_{K_{1}}\cap\{Z=1\}\hookrightarrow V_{K_{0}}\cap\{Z=1\}\colon(x,y)\mapsto(xy^{b},y)$
for some $b\in\mathbb{Z}$. This can be seen as a slight refinement of [17, Theorem 1.1], but it gives no constraint on $K_{0}$ if we suppose that $K_{1}$ is the unknot. The reason is that $V_{K_{1}}\cap\{Z=1\}=\{x=1\}\cup\{y=1\}$ if $K_{1}$ is the unknot (see [16, Example 3.13]) and $\{x=1\}\cup\{y=1\}$ is a subset of $V_{K_{0}}\cap\{Z=1\}$ for any knot $K_{0}$ (see [14, Proposition 5.6] and [16, Exercise 3.17]).
@@
Organization of paper. In Section 2, for a contact manifold $Y$ and its Legendrian submanifold $\Lambda$ satisfying certain conditions, we introduce the Chekanov-Eliashberg DGA of $(Y,\Lambda)$ with coefficients in $\mathbb{Z}[H_{2}(Y,\Lambda)]$. We then define a DGA map associated to an exact Lagrangian cobordism in ×$\mathbb{R}\times Y$. In Section 2.5, we will see how to apply this DGA map to a setting with a clean Lagrangian intersection as in Question 1.1. In Section 3, we combine the argument in Section 2.5 with the results of [4, 16] about the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$. Theorem 1.4 is proved in Section 3.1. Theorem 1.1 is proved by using Lemma 1.2 (Lemma 3.9) in Section 3.2. In Appendix A, by using Floer cohomology group over $\mathbb{F}_{2}[\lambda^{\pm}]$, we prove Proposition 2.7 which gives a homological constraint on $1$-dimensional clean intersections. This enhances some results in [17]. Nevertheless, the present proof is much simpler than that in [17].
Acknowledgements. The author would like to thank Kei Irie for making valuable comments. He would also like to thank Fumikazu Nagasato and Kouki Sato for fruitful discussion before this work and for giving some insight from knot theory. This work was supported by JSPS KAKENHI Grant Number JP23KJ1238.

### v2

# Topological constraints on clean Lagrangian intersections from $\mathbb{Q}$-valued augmentations

@@ Abstract
Let $K$ be the connected sum in $\mathbb{R}^{3}$ of an arbitrary knot and the $(2,q)$-torus knot for $q\neq\pm 1$, and consider its conormal bundle $L_{K}$ in $T^{*}\mathbb{R}^{3}$. We prove that there is no compactly supported Hamiltonian diffeomorphism $\varphi$ on $T^{*}\mathbb{R}^{3}$ such that $\varphi(L_{K})$ and the zero section $\mathbb{R}^{3}$ intersect cleanly along the unknot in $\mathbb{R}^{3}$. For the proof, we examine the knot DGA, whose homology was proved to be isomorphic to the knot contact homology by Ekholm, Etnyre, Ng and Sullivan, with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$. We find an algebraic constraint imposed on every $\mathbb{Q}$-valued augmentation of the knot DGA of $K$ when $q\geq 3$. This constraint involves rational roots of polynomials in $\mathbb{Q}[T]$, and we apply Hilbert’s irreducibility theorem to complete the proof. Using a similar idea, we also prove an analogous result replacing the $(2,q)$-torus knot with the figure-eight knot.
@@ 1 Introduction
It is a Lagrangian submanifold of $T^{*}\mathbb{R}^{3}$ and intersects $\mathbb{R}^{3}$ cleanly along the knot $K$. Let $\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ denote the set of Hamiltonian diffeomorphisms on $T^{*}\mathbb{R}^{3}$ associated to compactly supported Hamiltonians on $T^{*}\mathbb{R}^{3}$. Using these notations, we ask the following question, which was observed by the author in [16].
Question 1.1 (Question 1.1 of [16]).
@@
We remark that it is essential that $\varphi$ is a Hamiltonian diffeomorphism. Indeed, for any pair of knots $(K_{0},K_{1})$ in $\mathbb{R}^{3}$, there exists a $C^{\infty}$ isotopy $(\varphi^{t})_{t\in[0,1]}$ on $T^{*}\mathbb{R}^{3}$ with compact support such that $\varphi^{1}(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$. See [16, Proposition 4.1].
@@
As a special case, the answer to Question 1.1 when $K$ is the unknot is ‘no’. This can be deduced from the result by Ganatra-Pomerleano [9, Proposition 6.29] about clean Lagrangian intersections in a plumbing of two copies of the cotangent bundle of $S^{3}$ along the unknots (see [16, Section 1]). Smith-Wemyss [20, Proposition 4.10] and Asplund-Li [1, Theorem 1.3] also studies clean Lagrangian intersections in plumbings of two copies of $T^{*}S^{3}$ along the unknots. These results are concerned with Smith’s question on the ‘persistence (or rigidity) of unknottedness of Lagrangian intersections’ [9, Question 1.5].
@@
In a general case, under the assumption that $\varphi(L_{K})$ for $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ intersects $\mathbb{R}^{3}$ cleanly along a knot $K^{\prime}$ in $\mathbb{R}^{3}$, the author proved [16, Theorem 1.1]. It relates the augmentation variety in $(\mathbb{C}^{*})^{2}$ of $K$ with that of $K^{\prime}$. This result leads us to the same answer to Question 1.1 as above when $K$ is the unknot [16, Theorem 1.3]. We can also deduce some constraints on the knot type of $K^{\prime}$ when $K$ has certain non-trivial knot types [16, Theorem 1.4, 1.5].
Main results. In fact, the aforementioned result [16, Theorem 1.1] is useless to solve the following weaker version of Question 1.1. For the reason, see Remark 1.5 below.
Question 1.2.
@@
If $K$ is not the unknot, is there $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{n})$ such that $\varphi(L_{K})$ intersects $\mathbb{R}^{3}$ cleanly along the unknot in $\mathbb{R}^{3}$?
The goal of this paper is to give a negative answer to this question when $K$ has a certain knot type. As a notation, let $K_{1}\#K_{2}$ denote the connected sum of knots $K_{1}$ and $K_{2}$ in $\mathbb{R}^{3}$. Let $T_{(2,2m+1)}$ for $m\in\mathbb{Z}$ denote the $(2,2m+1)$-torus knot and let $4_{1}$ denote the figure-eight knot.
Theorem 1.1 (Theorem 3.15 and Theorem 3.19).
Let $K$ be $K^{\prime}\#T_{(2,2m+1)}$ or $K^{\prime}\#4_{1}$, where $K^{\prime}$ is an arbitrary knot in $\mathbb{R}^{3}$ and $m\in\mathbb{Z}\setminus\{0,-1\}$. Then, there is no $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K})$ intersects $\mathbb{R}^{3}$ cleanly along the unknot in $\mathbb{R}^{3}$.
@@
For the proof, the geometric arguments in this paper are similar to those in [16], which are outlined as follows: Let $K_{0}$ and $K_{1}$ be knots in $\mathbb{R}^{3}$. For $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$, suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ intersect cleanly along $K_{1}$. We use the results by Ekholm-Etnyre-Ng-Sullivan [6, 3] which show that the Chekanov-Eliashberg DGA of the unit conormal bundle of $K_{i}$ is stable tame isomorphic to the knot DGA of $K_{i}$ for $i=0,1$. The knot DGA, introduced by Ng [12], is defined combinatorially from a braid representation of a knot. Here, we refer to the one with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$ defined in [14, 3, 15]. We then apply a DGA map associated to an exact Lagrangian cobordism between the unit conormal bundles of $K_{0}$ and $K_{1}$. This is a part of the symplectic field theory introduced by Eliashberg-Givental-Hofer [8]. Consequently, we obtain a DGA map from the knot DGA of $K_{0}$ to that of $K_{1}$. From a simple description of the knot DGA of the unknot, we obtain the following result about augmentations of the knot DGA of $K_{0}$.
Proposition 1.2 (Proposition 3.5).
Let $\mathbb{k}$ be a field. Suppose that $K_{1}$ is the unknot. Then, for every $y,Z\in\mathbb{k}^{*}$ satisfying $y\notin\{Z^{n}\mid n\in\mathbb{Z}\}$, there exists an augmentation $\varepsilon$ of the knot DGA of $K_{0}$ to $\mathbb{k}$ such that $\varepsilon(\mu)=y$ and $\varepsilon(U)=Z$.
There are two new points in algebraic arguments in this paper. Let us explain them by focusing on the knot $K^{\prime}\#T_{(2,2m+1)}$ for $m\in\mathbb{Z}_{\geq 1}$.
The first point is that we find a constraint on augmentations of some knot DGAs in terms of roots of polynomials in $\mathbb{k}[T]$, where $\mathbb{k}$ is any field. Actually, we find a specific polynomial $P_{m}\in\mathbb{Z}[\mu,U][T]$ for every $m\in\mathbb{Z}_{\geq 1}$, whose degree in $T$ is $m+1$, with the following property (see Proposition 3.10): For any augmentation $\varepsilon$ of the knot DGA of $K^{\prime}\#T_{(2,2m+1)}$ to $\mathbb{k}$ such that $\varepsilon(\mu)=y$ and $\varepsilon(U)=Z$, the polynomial $\left.P_{m}\right|_{\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$. About the derivation of $P_{m}$, see Remark 1.3 below. We will focus on the case $\mathbb{k}=\mathbb{Q}$. It is worth noting that this constraint on augmentations is trivial if $\mathbb{k}$ is algebraically closed.
The second point is that we use a classical result in number theory. Indeed, Theorem 1.1 when $K=K^{\prime}\#T_{(2,2m+1)}$ for $m\in\mathbb{Z}_{\geq 1}$ is proved by a contradiction as follows: Suppose that $K_{0}=K^{\prime}\#T_{(2,2m+1)}$ and $K_{1}$ is the unknot. Then, Proposition 1.2 and the property of $P_{m}$ for $\mathbb{k}=\mathbb{Q}$ imply that $\left.P_{m}\right|_{\mu=2,U=Z}\in\mathbb{Q}[T]$ has a root in $\mathbb{Q}$ for every $Z\in\mathbb{Q}^{*}\setminus\{2^{\pm 1}\}$. On the other hand, by using Hilbert’s irreducibility theorem, we can show that there are infinitely many $Z\in\mathbb{Q}$ such that $\left.P_{m}\right|_{\mu=2,U=Z}\in\mathbb{Q}[T]$ does not have a root in $\mathbb{Q}$, which is a contradiction. (From the case of $m\geq 1$, we can deduce Theorem 1.1 when $K=K^{\prime}\#T_{(2,-2m-1)}$.)
The proof of Theorem 1.1 when $K=K^{\prime}\#4_{1}$ goes in a similar way. We explicitly give a polynomial $P\in\mathbb{Z}[\mu^{\pm},U^{\pm}][T]$, whose degree in $T$ is $3$, with the following property (see Proposition 3.17): For any augmentation $\varepsilon$ of the knot DGA of $K^{\prime}\#4_{1}$ to $\mathbb{k}$ such that $\varepsilon(\mu)=y$ and $\varepsilon(U)=Z$, the polynomial $\left.P\right|_{\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$. Using concrete calculations and the rational root theorem, we show that $\left.P\right|_{\mu=-1,U=-2}\in\mathbb{Q}[T]$ does not have a root in $\mathbb{Q}$. Combining with Proposition 1.2 for $\mathbb{k}=\mathbb{Q}$, Theorem 1.1 when $K=K^{\prime}\#4_{1}$ is deduced.
@@
The polynomials $P_{m}$ for $K^{\prime}\#T_{(2,2m+1)}$ and $P$ for $K^{\prime}\#4_{1}$ are found from concrete computations on some part of (the homology of) the knot DGA which comes from the component $T_{(2,2m+1)}$ or $4_{1}$ of the connected sum. Our computation relies on braid representations of $T_{(2,2m+1)}$ and $4_{1}$, which are rather simple. It might be possible to find a polynomial with a similar property as $P_{m}$ for more complicated knots, such as the $(p,q)$-torus knot when $3\leq p<|q|$.
We also provide a theorem which is a refinement of [16, Theorem 1.1]. For an oriented knot $K$ in $\mathbb{R}^{3}$, let $V_{K}$ denote the augmentation variety defined by [15, Definition 5.1] as a subset of $(\mathbb{C}^{*})^{3}$. Let $(x,y,Z)$ be the coordinate of $\mathbb{C}^{3}$.
@@
For oriented knots $K_{0}$ and $K_{1}$ in $\mathbb{R}^{3}$, suppose that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ intersects $\mathbb{R}^{3}$ cleanly along $K_{1}$. Then, there exist $a\in\{\pm 1\}$ and integers $b,c$ and $d$ for which we have an injective map
$V_{K_{1}}\hookrightarrow V_{K_{0}}\colon(x,y,Z)\mapsto(x^{a}y^{b}Z^{c},y^{a}Z^{d},Z).$
@@
Identifying $\{(x,y,Z)\in(\mathbb{C}^{*})^{3}\mid Z=1\}$ with $(\mathbb{C}^{*})^{2}$, we obtain $V_{K}\cap\{Z=1\}\subset(\mathbb{C}^{*})^{2}$, which is a version of augmentation variety used in [16]. This corresponds to the reduction of the coefficient ring of the knot DGAs from $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$ to $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ by setting $U=1$. It follows from Theorem 1.4 that there exists an injective map
$V_{K_{1}}\cap\{Z=1\}\hookrightarrow V_{K_{0}}\cap\{Z=1\}\colon(x,y)\mapsto(x^{a}y^{b},y^{a})$
for some $a\in\{\pm 1\}$ and $b\in\mathbb{Z}$. This is a slight refinement of [16, Theorem 1.1], but it gives no constraint on $K_{0}$ if we suppose that $K_{1}$ is the unknot. The reason is that $V_{K_{1}}\cap\{Z=1\}=\{x=1\}\cup\{y=1\}$ if $K_{1}$ is the unknot (see [15, Example 3.13]) and $\{x=1\}\cup\{y=1\}$ is a subset of $V_{K_{0}}\cap\{Z=1\}$ for every knot $K_{0}$ (see [13, Proposition 5.6] and [15, Exercise 3.17]).
@@
Organization of paper. In Section 2, for a contact manifold $Y$ and its Legendrian submanifold $\Lambda$ satisfying certain conditions, we introduce the Chekanov-Eliashberg DGA of $(Y,\Lambda)$ with coefficients in $\mathbb{Z}[H_{2}(Y,\Lambda)]$. We then define a DGA map associated to an exact Lagrangian cobordism in $\mathbb{R}\times Y$. In Section 2.5, we will see how to apply this DGA map to a setting with clean Lagrangian intersection as in Question 1.1. In Section 3, we combine the argument in Section 2.5 with the results of [3, 15] about the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$. Proposition 1.2 and Theorem 1.4 are proved in Section 3.1. Theorem 1.1 for $K=K^{\prime}\#T_{(2,2m+1)}$ with $m\neq 0,-1$ (resp. $K=K^{\prime}\#4_{1}$) is proved by examining the knot DGA of $K^{\prime}\#T_{(2,2m+1)}$ in Section 3.2 (resp. $K^{\prime}\#4_{1}$ in Section 3.3). In Appendix A, by using Lagrangian Floer cohomology group over $\mathbb{F}_{2}[\lambda^{\pm}]$, we prove Proposition 2.7 which gives a homological constraint on $1$-dimensional clean intersections. This strengthen a result in [16]. Nevertheless, the present proof is much simpler than that of [16].
Acknowledgements. The author would like to thank Kei Irie for making valuable comments. He would also like to thank Fumikazu Nagasato and Kouki Sato for the discussion before this work and giving some insight from knot theory. He would also like to thank Lenhard Ng for the discussion during CBMS summer school 2025 and suggestions for improving the main result. The author appreciates the referee’s valuable comments. This work was supported by JSPS KAKENHI Grant Number JP23KJ1238. The author is currently supported by JSPS KAKENHI Grant Number JP25KJ0270.